# A Survey of Reinforcement Learning For Economics

### 15 Mar 2026 v2

@@ 8.4 Financial Order Execution
Nevmyvaka et al. (2006) applied tabular Q-learning to execution on real limit order book data from NASDAQ stocks.123123123The dataset covers 500 trading days of millisecond-level limit order book snapshots for AMZN, QCOM, and NVDA. The state space contains approximately 10,000 states; the horizon is $T=60$ seconds with discount $\gamma=1$. The state at time $t$ is $s_{t}=(t,q_{t},\psi_{t},\Delta_{t})$, where $t\in\{1,\ldots,T\}$ is time remaining, $q_{t}\in\{0,1,\ldots,Q\}$ is inventory remaining, $\psi_{t}$ is the discretized bid-ask spread, and $\Delta_{t}$ is the discretized signed volume imbalance.124124124Signed volume imbalance is the difference between buy and sell order volume near the best prices, normalized by total volume; positive imbalance typically predicts short-term price increases. Actions $a_{t}\in\{0,\delta,2\delta,\ldots,q_{t}\}$ specify shares to execute in the current interval. The reward is the negative per-period slippage contribution:
@@
where $p_{0}$ is the mid-quote at order arrival and $p_{t}^{\text{exec}}$ is the average execution price. Total implementation shortfall is $IS=-\sum_{t=1}^{T}r_{t}/Q$.125125125Implementation shortfall measures the total cost of executing a trade relative to the benchmark mid-quote at arrival. It captures market impact, timing cost, and opportunity cost of unexecuted shares. Because the objective is cost minimization, the Q-learning update uses $\min$ over next-period actions:
@@ 8.5 Supply Chain Inventory Management
Gijsbrechts et al. (2022) conducted a systematic evaluation of deep RL against classical base-stock policies across lost sales, dual sourcing, and multi-echelon configurations. The classical benchmark is the echelon base-stock policy, in which each stage $k$ maintains an echelon inventory position and orders to bring this position to a target level $S_{k}$.126126126Formally, the echelon inventory position at stage $k$ equals on-hand inventory at $k$ plus all inventory at downstream stages $1,\ldots,k-1$ plus in-transit inventory, minus backorders. The echelon base-stock policy of Clark and Scarf (1960) is optimal for serial systems with linear costs and backorders. For single-echelon systems with backorders, the optimal base-stock level is given by the newsvendor critical fractile $S^{*}=F_{D}^{-1}(b/(b+h))$. Table 15 summarizes results from their multi-echelon experiments.
@@ 6 $\sim 10^{8}$ Failed
Where classical solutions exist, RL underperforms: the base-stock policy remains highly competitive when properly calibrated, and DRL required millions of training transitions to approach performance levels achievable through closed-form calculation. RL struggles particularly with multi-echelon coupling, where upstream orders affect downstream costs many periods later.127127127Credit assignment refers to the difficulty of determining which past actions caused a delayed reward or cost. In multi-echelon systems, an upstream ordering decision today may not affect customer-facing costs for many periods, making it difficult for RL to learn the causal connection. The value proposition for RL lies in problems where analytical solutions are unavailable: non-stationary demand, complex operational constraints, or cost structures that do not admit tractable decomposition. Even in these settings, successful deployment remains rare and requires extensive simulation infrastructure, domain expertise, and careful calibration against classical baselines.
@@ 8.6 Real-Time Bidding
where $\mathrm{cvr}_{t}\in\{0,1\}$ is the conversion indicator. Budget evolves as $B_{t+1}=B_{t}-\mathbbm{1}\{b_{t}\geq c_{t}\}\cdot c_{t}$, and the episode terminates when $B_{t}\leq 0$ or t=0t=0.128128128The state space is discretized into budget bins and time bins. Wu et al. (2018) augment the state with bid landscape features derived from historical clearing price distributions, giving the agent information about current market competitiveness.
@@ 8.7 Simulation Study: Bus Engine Replacement
I extend the single-engine problem to a fleet of $N$ engines with a capacity constraint limiting replacements per period. The state $s=(m_{1},\ldots,m_{N})$ records discretized mileage for each engine. Actions are subsets of engines to replace, subject to the capacity constraint. The per-period cost is $c(s,a)=\alpha\sum_{i}m_{i}+\beta|a|$, where $\alpha$ is the operating cost per unit mileage and $\beta$ is the replacement cost.129129129The mileage-dependent operating cost $\alpha\sum_{i}m_{i}$ follows Rust’s original specification $c(x,\theta_{1})=\theta_{11}x$, creating a non-trivial threshold replacement policy. The fleet extension uses deterministic mileage increments to isolate the combinatorial scaling challenge that arises from the joint state of multiple engines. Mileage evolves deterministically; replaced engines reset to $m=0$; others increment by one bin.130130130With $M=6$ mileage bins, the state space is $6^{N}$: 1,296 states at $N=4$, 7,776 at $N=5$, 46,656 at $N=6$. Value iteration is feasible for $N\leq 5$.
Table 17 compares dynamic programming, DQN, and heuristic baselines across fleet sizes.
Figure 15: Bus engine replacement benchmark. Left: computation time vs. fleet size (log scale). Right: discounted return vs. fleet size for DP, DQN, and heuristic baselines.
Table 17: Bus engine replacement results across fleet sizes.
NN $|\mathcal{S}|$ DP Time DP Return DQN Return Thresh(3) Never Replace
1 6 $0.06$s 55.1-55.1 $5.1\pm 0.1 55$.7-55.7 94.0-94.0
2 36 $0.06$s 110.3-110.3 $10.0\pm 0.3 11$1.4-111.4 187.7-187.7
3 216 $0.59$s 165.3-165.3 $165.0\pm 0.1$ 168.3-168.3 282.1-282.1
4 1,296 $6.93$s 220.0-220.0 $20.8\pm 0.1 22$3.8-223.8 376.0-376.0
5 7,776 $81.79$s 275.7-275.7 $276.2\pm 0.1$ 279.5-279.5 469.7-469.7
6 46,656 $31.7\pm 1.1 33$6.6-336.6 562.8-562.8
@@ 9 Structural Estimation with Reinforcement Learning
Several recent papers have used RL training loops, namely Q-learning, temporal-difference learning, policy gradient, and actor-critic methods, to solve structural economic models at scales where conventional dynamic programming fails.131131131I exclude papers that use neural network function approximation without an RL training mechanism. Inverse reinforcement learning is treated in the sister survey (Rust and Rawat, 2026). Throughout this chapter I adopt a unified notation. An MDP is a tuple $(\mathcal{S},\mathcal{A},P,r,\gamma)$ where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P(s^{\prime}|s,a)$ is the transition kernel, $r(s,a)$ is the per-period reward, and $\gamma\in[0,1)$ is the discount factor.132132132Several of the papers reviewed here use $\beta$ for the discount factor, following economics convention. I translate all results to $\gamma$ for consistency with the RL literature and the rest of this survey. A policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ maps states to distributions over actions. The value function under $\pi$ is $V^{\pi}(s)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\mid s_{0}=s\right]$, and the action-value function is $Q^{\pi}(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[V^{\pi}(s^{\prime})]$. The optimal value function satisfies $V^{*}(s)=\max_{a\in\mathcal{A}}Q^{*}(s,a)$.
@@ 9.0.1 Adusumilli and Eckardt (2022): TD Learning for CCP Estimation
The first is the linear semi-gradient method.133133133The method is called “semi-gradient” because it computes the gradient of only the prediction $\phi(a,s)^{\top}w$ with respect to $w$, not the full TD error including the bootstrap target $\phi(a^{\prime},s^{\prime})^{\top}w$. This avoids differentiating through the target but sacrifices guaranteed convergence in some settings. Approximate $h(a,s)\approx\phi(a,s)^{\top}w$ where $\phi:\mathcal{A}\times\mathcal{S}\to\mathbb{R}^{p}$ is a vector of basis functions (e.g., polynomials in state variables) and $w\in\mathbb{R}^{p}$ are the weights to be estimated. The TD(0) fixed-point equation is
@@
and then fits $\hat{h}^{(k)}$ by regressing $Y_{it}^{(k)}$ on $(a_{it},s_{it})$ using any machine learning method, including LASSO, random forests, or neural networks.134134134LASSO (Tibshirani, 1996) adds an $\ell_{1}$ penalty to a regression loss, shrinking many coefficients to zero for sparse solutions; random forests average many decision trees fit to random subsamples and feature subsets for nonparametric estimation; neural networks compose affine transformations with elementwise nonlinearities across multiple layers. This is the first DDC estimator compatible with arbitrary ML prediction methods, enabling application to very high-dimensional state spaces.
With $\hat{h}$ and $\hat{g}$ in hand, structural parameters are recovered by pseudo-maximum likelihood estimation (PMLE). For continuous state spaces, Adusumilli et al. (2022) derive a locally robust correction to the PMLE criterion that accounts for the nonparametric first-stage estimation of value terms, restoring $\sqrt{n}$-convergence of $\hat{\theta}$.135135135An estimator has $\sqrt{n}$-convergence if its error shrinks at rate $n^{-1/2}$ where $n$ is the sample size. This is the standard parametric rate; slower rates (e.g., $n^{-1/4}$) indicate efficiency loss from nonparametric first stages. The PMLE score is $m(a,s;\theta,h,g)=\partial_{\theta}\ln\pi(a,s;\theta,h,g)$, where $\pi$ is the logit choice probability with continuation value $V(a,s)=h(a,s)^{\top}\theta+g(a,s)$. The naive estimator solves $\mathbb{E}_{n}[m(a,s;\theta,\hat{h},\hat{g})]=0$, but with continuous states this moment condition is not orthogonal to the first-stage estimates and converges slower than $\sqrt{n}$. The locally robust moment adds a debiasing correction:
@@
where $\lambda(a,s;\theta)$ solves a backward recursion that propagates the influence of estimation error through the dynamic structure.136136136The term in braces is the temporal-difference error of the continuation value $V$. The adjoint $\lambda$ weights this TD error by its marginal impact on the PMLE score. See Online Appendix B.3 of Adusumilli et al. (2022) for the derivation. The corrected estimator $\hat{\theta}_{LR}$ solves $\mathbb{E}_{n}[\zeta_{n}]=0$ and is computationally no harder than the naive PMLE, since the correction is constant in $\theta$.
@@
Under regularity conditions, the linear semi-gradient estimator $\hat{h}$ satisfies $\|\hat{h}-h\|_{2}=O_{P}(n^{-1/2}(T-1)^{-1/2})$, where $\|\cdot\|_{2}$ denotes the $L^{2}(P)$ norm.137137137The $L^{2}(P)$ norm is $\|f\|_{2}=(\int f(x)^{2}\,dP(x))^{1/2}$, measuring average squared deviation under the probability measure $P$. This is the natural norm for mean-squared-error analysis.
@@ 9.0.2 Hu and Yang (2025): Policy Gradient for DDC Estimation
The outer loop is SMM.138138138SMM (Gourieroux et al., 1993) estimates structural parameters by matching moments from simulated data to moments from observed data, avoiding direct evaluation of the likelihood function. Define a vector of data moments $\mathbf{M}_{d}$ computed from the observed panel. For candidate structural parameters $\theta$ and transition parameters $\xi$, simulate the model to produce simulated moments $\mathbf{M}_{s}(\theta,\xi)$. The estimator minimizes
@@
The inner loop parametrizes the choice probability directly as a logistic function of state variables. For a binary choice $J_{t}\in\{0,1\}$, the general form is $\Pr(J_{t}=1\mid\boldsymbol{X}_{t};\boldsymbol{\gamma}(\theta))=\text{logistic}(\boldsymbol{X}_{t}\boldsymbol{\gamma}(\theta))$, where $\boldsymbol{\gamma}(\theta)$ are policy parameters that depend on the structural parameters.139139139The linear index can be replaced by higher-order terms of $\boldsymbol{X}_{t}$ or deep neural networks; the method requires only that the gradient $\nabla_{\boldsymbol{\gamma}}\log\pi_{\boldsymbol{\gamma}}$ has a closed form. In their application to a Rust bus engine model with unobserved bus condition $S_{t}^{*}$, this takes the form
@@ 9.1.2 Hollenbeck (2019): TD Learning for Merger Analysis with Innovation
The value function update for firm $i$ is142142142This stochastic approximation update, introduced by Pakes and McGuire (1994) for dynamic oligopoly computation, is closely related to temporal-difference learning in the RL literature. The original algorithm uses visit-count averaging ($\alpha_{k}=1/k$ where $k$ counts visits to each state) rather than a fixed learning rate.
@@
where $\Pi_{i}(\Omega,\mathbf{a})$ denotes firm $i$’s per-period profit given industry state $\Omega$ and the joint action profile $\mathbf{a}$ (investment, entry/exit, merger decisions),143143143I use $\Pi_{i}$ for firm $i$’s per-period profit to avoid confusion with policy $\pi$. $\gamma$ is the discount factor, and $\Omega^{\prime}$ is the realized next-period state. The algorithm uses $\varepsilon$-decreasing exploration to prevent convergence to locally suboptimal equilibria.
The equilibrium computation follows the Pakes-McGuire iterative scheme.144144144The Pakes-McGuire algorithm (Pakes and McGuire, 1994) computes Markov perfect equilibria by iterating: (1) given current value functions, compute best-response strategies; (2) given strategies, update value functions via simulation. Convergence is not guaranteed but works well in practice. The algorithm simulates long industry histories, updates each firm’s value function via (80), re-derives best-response strategies from updated values, and repeats.
@@
Two related papers merit brief mention. Lomys and Magnolfi (2024) develop structural estimation methods for strategic settings where agents use learning algorithms (specifically regret-minimizing rules) rather than playing a fixed equilibrium. They impose an “asymptotic no-regret” condition as a minimal rationality requirement and derive identification results for payoff parameters. Covarrubias (2022) uses deep RL to study oligopolistic pricing in a New Keynesian framework, representing firms’ pricing policies as neural networks $\pi_{\phi}(a|s)$. The method uncovers multiple equilibria ranging from competitive to collusive pricing.145145145The most prominent example of algorithmic collusion is Calvano et al. (2020), who showed that independent Q-learning agents in a repeated Bertrand pricing game learn to sustain supra-competitive prices and punish deviators without explicit communication. This result and its implications for competition policy are treated in the companion thesis chapter (Rawat, 2026).

### 17 Mar 2026 v3

@@ 7.4 Financial Order Execution
Nevmyvaka et al. (2006) applied tabular Q-learning to execution on real limit order book data from NASDAQ stocks.129129129The dataset covers 500 trading days of millisecond-level limit order book snapshots for AMZN, QCOM, and NVDA. The state space contains approximately 10,000 states; the horizon is $T=60$ seconds with discount $\gamma=1$. The state at time $t$ is $s_{t}=(t,q_{t},\psi_{t},\Delta_{t})$, where $t\in\{1,\ldots,T\}$ is time remaining, $q_{t}\in\{0,1,\ldots,Q\}$ is inventory remaining, $\psi_{t}$ is the discretized bid-ask spread, and $\Delta_{t}$ is the discretized signed volume imbalance.130130130Signed volume imbalance is the difference between buy and sell order volume near the best prices, normalized by total volume; positive imbalance typically predicts short-term price increases. Actions $a_{t}\in\{0,\delta,2\delta,\ldots,q_{t}\}$ specify shares to execute in the current interval. The reward is the negative per-period slippage contribution:
@@
where $p_{0}$ is the mid-quote at order arrival and $p_{t}^{\text{exec}}$ is the average execution price. Total implementation shortfall is $IS=-\sum_{t=1}^{T}r_{t}/Q$.131131131Implementation shortfall measures the total cost of executing a trade relative to the benchmark mid-quote at arrival. It captures market impact, timing cost, and opportunity cost of unexecuted shares. Because the objective is cost minimization, the Q-learning update uses $\min$ over next-period actions:
@@ 7.5 Supply Chain Inventory Management
Gijsbrechts et al. (2022) conducted a systematic evaluation of deep RL against classical base-stock policies across lost sales, dual sourcing, and multi-echelon configurations. The classical benchmark is the echelon base-stock policy, in which each stage $k$ maintains an echelon inventory position and orders to bring this position to a target level $S_{k}$.132132132Formally, the echelon inventory position at stage $k$ equals on-hand inventory at $k$ plus all inventory at downstream stages $1,\ldots,k-1$ plus in-transit inventory, minus backorders. The echelon base-stock policy of Clark and Scarf (1960) is optimal for serial systems with linear costs and backorders. For single-echelon systems with backorders, the optimal base-stock level is given by the newsvendor critical fractile $S^{*}=F_{D}^{-1}(b/(b+h))$. Table 13 summarizes results from their multi-echelon experiments.
@@ 6 $\sim 10^{8}$ Failed
Where classical solutions exist, RL underperforms: the base-stock policy remains highly competitive when properly calibrated, and DRL required millions of training transitions to approach performance levels achievable through closed-form calculation. RL struggles particularly with multi-echelon coupling, where upstream orders affect downstream costs many periods later.133133133Credit assignment refers to the difficulty of determining which past actions caused a delayed reward or cost. In multi-echelon systems, an upstream ordering decision today may not affect customer-facing costs for many periods, making it difficult for RL to learn the causal connection. The value proposition for RL lies in problems where analytical solutions are unavailable: non-stationary demand, complex operational constraints, or cost structures that do not admit tractable decomposition. Even in these settings, successful deployment remains rare and requires extensive simulation infrastructure, domain expertise, and careful calibration against classical baselines.
@@ 7.6 Real-Time Bidding
where $\mathrm{cvr}_{t}\in\{0,1\}$ is the conversion indicator. Budget evolves as $B_{t+1}=B_{t}-\mathbbm{1}\{b_{t}\geq c_{t}\}\cdot c_{t}$, and the episode terminates when $B_{t}\leq 0$ or t=0t=0.134134134The state space is discretized into budget bins and time bins. Wu et al. (2018) augment the state with bid landscape features derived from historical clearing price distributions, giving the agent information about current market competitiveness.
@@ 7.7 Simulation Study: Bus Engine Replacement
I extend the single-engine problem to a fleet of $N$ engines with a capacity constraint limiting replacements per period. The state $s=(m_{1},\ldots,m_{N})$ records discretized mileage for each engine. Actions are subsets of engines to replace, subject to the capacity constraint. The per-period cost is $c(s,a)=\alpha\sum_{i}m_{i}+\beta|a|$, where $\alpha$ is the operating cost per unit mileage and $\beta$ is the replacement cost.135135135The mileage-dependent operating cost $\alpha\sum_{i}m_{i}$ follows Rust’s original specification $c(x,\theta_{1})=\theta_{11}x$, creating a non-trivial threshold replacement policy. The fleet extension uses deterministic mileage increments to isolate the combinatorial scaling challenge that arises from the joint state of multiple engines. Mileage evolves deterministically; replaced engines reset to $m=0$; others increment by one bin.136136136With $M=6$ mileage bins, the state space is $6^{N}$: 1,296 states at $N=4$, 7,776 at $N=5$, 46,656 at $N=6$. Value iteration is feasible for $N\leq 5$.
Figure 13 compares dynamic programming, DQN, and heuristic baselines across fleet sizes.
Figure 13: Bus engine replacement benchmark. Left: computation time vs. fleet size (log scale). Right: discounted return vs. fleet size for DP, DQN, and heuristic baselines. At $N=6$ (46,656 states), DP is infeasible (no data point).
@@ 8 Structural Estimation with Reinforcement Learning
Several recent papers have used RL training loops,137137137Throughout this chapter, the RL training loop runs entirely inside the econometrician’s computational model; the agent never interacts with real economic agents or markets but serves as a numerical method for solving the Bellman equation within a structural estimation procedure (Section 2). There is no execution phase in the usual sense. namely Q-learning, temporal-difference learning, policy gradient, and actor-critic methods, to solve structural economic models at scales where conventional dynamic programming fails.138138138I exclude papers that use neural network function approximation without an RL training mechanism. Inverse reinforcement learning is treated in the sister survey (Rust and Rawat, 2026). Throughout this chapter I adopt a unified notation. An MDP is a tuple $(\mathcal{S},\mathcal{A},P,r,\gamma)$ where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P(s^{\prime}|s,a)$ is the transition kernel, $r(s,a)$ is the per-period reward, and $\gamma\in[0,1)$ is the discount factor.139139139Several of the papers reviewed here use $\beta$ for the discount factor, following economics convention. I translate all results to $\gamma$ for consistency with the RL literature and the rest of this survey. A policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ maps states to distributions over actions. The value function under $\pi$ is $V^{\pi}(s)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\mid s_{0}=s\right]$, and the action-value function is $Q^{\pi}(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[V^{\pi}(s^{\prime})]$. The optimal value function satisfies $V^{*}(s)=\max_{a\in\mathcal{A}}Q^{*}(s,a)$.
@@ 8.0.1 Adusumilli and Eckardt (2022): TD Learning for CCP Estimation
The first is the linear semi-gradient method.140140140The method is called “semi-gradient” because it computes the gradient of only the prediction $\phi(a,s)^{\top}w$ with respect to $w$, not the full TD error including the bootstrap target $\phi(a^{\prime},s^{\prime})^{\top}w$. This avoids differentiating through the target but sacrifices guaranteed convergence in some settings. Approximate $h(a,s)\approx\phi(a,s)^{\top}w$ where $\phi:\mathcal{A}\times\mathcal{S}\to\mathbb{R}^{p}$ is a vector of basis functions (e.g., polynomials in state variables) and $w\in\mathbb{R}^{p}$ are the weights to be estimated. The TD(0) fixed-point equation is
@@
and then fits $\hat{h}^{(k)}$ by regressing $Y_{it}^{(k)}$ on $(a_{it},s_{it})$ using any machine learning method, including LASSO, random forests, or neural networks.141141141LASSO (Tibshirani, 1996) adds an $\ell_{1}$ penalty to a regression loss, shrinking many coefficients to zero for sparse solutions; random forests average many decision trees fit to random subsamples and feature subsets for nonparametric estimation; neural networks compose affine transformations with elementwise nonlinearities across multiple layers. This is the first DDC estimator compatible with arbitrary ML prediction methods, enabling application to very high-dimensional state spaces.
With $\hat{h}$ and $\hat{g}$ in hand, structural parameters are recovered by pseudo-maximum likelihood estimation (PMLE). For continuous state spaces, Adusumilli et al. (2022) derive a locally robust correction to the PMLE criterion that accounts for the nonparametric first-stage estimation of value terms, restoring $\sqrt{n}$-convergence of $\hat{\theta}$.142142142An estimator has $\sqrt{n}$-convergence if its error shrinks at rate $n^{-1/2}$ where $n$ is the sample size. This is the standard parametric rate; slower rates (e.g., $n^{-1/4}$) indicate efficiency loss from nonparametric first stages. The PMLE score is $m(a,s;\theta,h,g)=\partial_{\theta}\ln\pi(a,s;\theta,h,g)$, where $\pi$ is the logit choice probability with continuation value $V(a,s)=h(a,s)^{\top}\theta+g(a,s)$. The naive estimator solves $\mathbb{E}_{n}[m(a,s;\theta,\hat{h},\hat{g})]=0$, but with continuous states this moment condition is not orthogonal to the first-stage estimates and converges slower than $\sqrt{n}$. The locally robust moment adds a debiasing correction:
@@
where $\lambda(a,s;\theta)$ solves a backward recursion that propagates the influence of estimation error through the dynamic structure.143143143The term in braces is the temporal-difference error of the continuation value $V$. The adjoint $\lambda$ weights this TD error by its marginal impact on the PMLE score. See Online Appendix B.3 of Adusumilli et al. (2022) for the derivation. The corrected estimator $\hat{\theta}_{LR}$ solves $\mathbb{E}_{n}[\zeta_{n}]=0$ and is computationally no harder than the naive PMLE, since the correction is constant in $\theta$.
@@
Under regularity conditions, the linear semi-gradient estimator $\hat{h}$ satisfies $\|\hat{h}-h\|_{2}=O_{P}(n^{-1/2}(T-1)^{-1/2})$, where $\|\cdot\|_{2}$ denotes the $L^{2}(P)$ norm.144144144The $L^{2}(P)$ norm is $\|f\|_{2}=(\int f(x)^{2}\,dP(x))^{1/2}$, measuring average squared deviation under the probability measure $P$. This is the natural norm for mean-squared-error analysis.
@@ 8.0.2 Hu and Yang (2025): Policy Gradient for DDC Estimation
The outer loop is SMM.145145145SMM (Gourieroux et al., 1993) estimates structural parameters by matching moments from simulated data to moments from observed data, avoiding direct evaluation of the likelihood function. Define a vector of data moments $\mathbf{M}_{d}$ computed from the observed panel. For candidate structural parameters $\theta$ and transition parameters $\xi$, simulate the model to produce simulated moments $\mathbf{M}_{s}(\theta,\xi)$. The estimator minimizes
@@
The inner loop parametrizes the choice probability directly as a logistic function of state variables. For a binary choice $J_{t}\in\{0,1\}$, the general form is $\Pr(J_{t}=1\mid\boldsymbol{X}_{t};\boldsymbol{\gamma}(\theta))=\text{logistic}(\boldsymbol{X}_{t}\boldsymbol{\gamma}(\theta))$, where $\boldsymbol{\gamma}(\theta)$ are policy parameters that depend on the structural parameters.146146146The linear index can be replaced by higher-order terms of $\boldsymbol{X}_{t}$ or deep neural networks; the method requires only that the gradient $\nabla_{\boldsymbol{\gamma}}\log\pi_{\boldsymbol{\gamma}}$ has a closed form. In their application to a Rust bus engine model with unobserved bus condition $S_{t}^{*}$, this takes the form
@@ 8.1.2 Hollenbeck (2019): TD Learning for Merger Analysis with Innovation
The value function update for firm $i$ is149149149This stochastic approximation update, introduced by Pakes and McGuire (1994) for dynamic oligopoly computation, is closely related to temporal-difference learning in the RL literature. The original algorithm uses visit-count averaging ($\alpha_{k}=1/k$ where $k$ counts visits to each state) rather than a fixed learning rate.
@@
where $\Pi_{i}(\Omega,\mathbf{a})$ denotes firm $i$’s per-period profit given industry state $\Omega$ and the joint action profile $\mathbf{a}$ (investment, entry/exit, merger decisions),150150150I use $\Pi_{i}$ for firm $i$’s per-period profit to avoid confusion with policy $\pi$. $\gamma$ is the discount factor, and $\Omega^{\prime}$ is the realized next-period state. The algorithm uses $\varepsilon$-decreasing exploration to prevent convergence to locally suboptimal equilibria.
The equilibrium computation follows the Pakes-McGuire iterative scheme.151151151The Pakes-McGuire algorithm (Pakes and McGuire, 1994) computes Markov perfect equilibria by iterating: (1) given current value functions, compute best-response strategies; (2) given strategies, update value functions via simulation. Convergence is not guaranteed but works well in practice. The algorithm simulates long industry histories, updates each firm’s value function via (78), re-derives best-response strategies from updated values, and repeats.
@@
Two related papers merit brief mention. Lomys and Magnolfi (2024) develop structural estimation methods for strategic settings where agents use learning algorithms (specifically regret-minimizing rules) rather than playing a fixed equilibrium. They impose an “asymptotic no-regret” condition as a minimal rationality requirement and derive identification results for payoff parameters. Covarrubias (2022) uses deep RL to study oligopolistic pricing in a New Keynesian framework, representing firms’ pricing policies as neural networks $\pi_{\phi}(a|s)$. The method uncovers multiple equilibria ranging from competitive to collusive pricing.152152152The most prominent example of algorithmic collusion is Calvano et al. (2020), who showed that independent Q-learning agents in a repeated Bertrand pricing game learn to sustain supra-competitive prices and punish deviators without explicit communication. This result and its implications for competition policy are treated in the companion thesis chapter (Rawat, 2026).